Circular No. 8109

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003cv Further to IAUC 8103, M. Hamuy reports that a spectrum (range 380–

930 nm) of SN 2003cv (cf. IAUC 8104), obtained by J. Maza and L. Huerta on Apr. 6.15 UT with the Dupont telescope, shows it to be a peculiar type-II supernova. Hα and Hβ lines exhibit very broad P-Cyg profiles; their minima yield expansion velocities of −14000 and −11300 km/s (based on a redshift of 0.028, derived for the host galaxy from superimposed narrow Balmer lines), which are unusually high for type-II events. The spectrum has a broad trough centered at 589.1 nm, probably due to Na I D 589.3-nm, and a P-Cyg profile due to the near-infrared Ca II triplet. This spectrum is very similar to that of the H-rich ‘hypernova’ SN 2003bg (IAUC 8084, 8088) taken on Apr. 4.04 by Maza and Huerta, and this suggests that SN 2003cv is at least least one month old.

IGR J16358–4726 C. Kouveliotou, Marshall Space Flight Center (MSFC) and Universities

Space Research Association (USRA); S. Patel, National Research Council; A. Tennant, MSFC; P. Woods and M. Finger, USRA; and S. Wachter, SIRTF Science Center, California Institute of Technology, report that IGR J16358–4726 (cf. IAUC 8097) was observed for 25700 s serendipitously with the Chandra X-ray Observatory during a scheduled observation of SGR 1627–41 starting on Mar. 24.17 UT. The source was ≈ 9′.7 off axis. A fit to the off-axis image was found to be consistent with the point-spread function and put the transient at α = 16^{h}35^{m}53^{s}.8, δ = −47^{o}25′41′′.1 (equinox 2000.0; error radius 0′′.6). The spectrum of the source is hard and highly absorbed, consistent with a power law of index 0.5(1) plus a line feature at 6.4 keV and a column density of N_{H} = 3.3 × 10^{23} cm^{−2}. The source flux (unabsorbed) is 1.7 × 10^{−10} (2–10 keV). Pulsations with a period of 5850(50) s are directly visible in the x-ray light curve with an energy-dependent amplitude; the pulse fraction between 2 and 10 keV is 63(6) percent, peak-to-peak. It is unclear whether these pulsations reflect a spin or orbital period, or a highcoherence quasiperiodic oscillation. A search for an infrared counterpart in the 2MASS All Sky Catalog reveals a relatively bright source 1′′.7 from the Chandra position of IGR J16358–4726 (2MASS J16355369–4725398, with J = 15.41, H = 13.44, K = 12.59). Further observations are encouraged.

SUPERNOVA 2003cz Corrigendum. On IAUC 8106, line 5, for Apr. 2.14 read Apr. 2.21

2003 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green